OpenStudy (anonymous):

What is the solution to the equation? 3q + 5 = –2q + 20 A. q = 3 B. q = 5 C. q = 15 D. q = 25 @iGreen

OpenStudy (igreen):

First we subtract 5 to both sides. What's 20 - 5?

OpenStudy (anonymous):

15

OpenStudy (igreen):

Yes, now we have: 3q = -2q + 15 Now we add 2q to both sides, what's 3q + 2q?

OpenStudy (anonymous):

5q

OpenStudy (anonymous):

3q + 5 = -q + 20 3q + 5) + -5 + 2q) = (-2q+ 20) + (-5 + 2g) (3q + 5) + (2q - 5) = (-2g + 20) + (2q - 5) = ?

OpenStudy (igreen):

Yes..so that gives us: 5q = 15 Now divide 5 to both sides, what do you get?

OpenStudy (anonymous):

1q=3?

OpenStudy (igreen):

Yep, you got it..3 is your answer.

OpenStudy (anonymous):

thanks green